# LAPTH-015/24 Spherical symmetric solutions of

### v2

@@
[…]
${}^{a}$ LAPTh, Université Savoie Mont Blanc, CNRS, F-74940 Annecy, France
${}^{b}$ LPTh, Department of Physics, University of Jijel,
@@ 1 Introduction
In the pursuit of understanding the universe’s accelerated expansion [2, 3], numerous alternatives have been proposed, both within the framework of general relativity and beyond. One such alternative involves introducing a new matter component termed dark energy, characterized by negative pressure. While the cosmological constant $\Lambda$, originally introduced by Einstein, remains the primary candidate for dark energy due to its alignment with observational data, challenges such as the fine-tuning and coincidence problems persist [4, 5, 6, 7]. On the other hand, a different path follows the idea that general relativity fails to describe the universe at larger scales and hence it should be modified. These modifications often propose alterations to fundamental principles of physics, seeking to reconcile observational evidence with theoretical predictions [8, 9]. Consequently, alternative models like scalar tensor theories [10, 11], Einstein-æther theory [12, 13], and higher derivative theories of gravity like $f(R)$ theories [14, 15] have been explored. Other kinds of modified theories are the modified teleparallel equivalent of general relativity, $f(T)$ gravity where the Ricci scalar is replaced by the torsion scalar, and the theory grounded in Weyl geometry, termed $f(Q)$ gravity [16]. The latter describes gravitational effects in terms of non-metricity, providing a broader framework than traditional Riemannian geometry. These theories offer different perspectives on gravity that may lead to new insights into cosmological and astrophysical phenomena.
@@ 2 Static spherically symmetric solutions with Maxwell source
Case 2: $d\neq 0,\,c\neq 0.$ The metric has the Lorentzian signature in three subcases: 2a) $c>0,\,d>0$, with $r$ real; 2b) $c<0,\,d>0$, with $r^{2}>\mu^{2}$; 2c) $c>0,\,d<0$, with $r^{2}<\mu^{2}$. In these three cases, the modulus of $c$ can be rescaled to $1$ by a rescaling of the time coordinate, so we can fix $c=\epsilon,\,d=\eta/\mu^{2}$, where $\epsilon=\pm 1$, $\eta=\pm 1$, and $\mu$ is a constant with the dimension of a length. The generic solution for these three cases is
@@
$+k_{2}\left[\displaystyle\frac{\sqrt{\eta r^{2}+\epsilon\mu^{2}}}{r}\,\psi\left(\displaystyle\frac{r}{\mu}\right)-1\right]-\frac{\Lambda r^{2}}{3},$
@@
$a)\psi(x)=$ ${\rm arsinh}\,{x},\quad\epsilon=+1,\;\eta=+1,$ (2.18)
$b)\psi(x)=$ ${\rm arcosh}\,{x},\quad\epsilon=-1,\;\eta=+1$ $\quad(r^{2}>\mu^{2}),$ (2.19)
@@
Expanding these solutions in Laurent series for the variable $r/\mu$, we recover the solutions given by Barnes [22].
@@ 3.1 Case 1
If $\lambda\neq 0$, the Ricci scalar [17]
@@
diverges on the sphere at spacelike infinity $r\to\infty$, which is a naked singularity of this spacetime. Furthermore, the coordinate transformation $r=x^{-1}$ leads to the asymptotic form of the metric
@@
showing that ”spacelike infinity” is actually at finite geodesic distance for all geodesics if $\lambda>0$, and for spacelike radial geodesics if $\lambda<0$. So the case $\lambda\neq 0$ is unphysical. When $\lambda=0$ the solution reduces to the Reissner-Nordström-(anti-)de Sitter solution of general relativity.
@@ 3.2 Case 2
$\mu^{2}R$ $=$ $-6\eta\left[\displaystyle\frac{\epsilon q^{2}}{r^{2}}+\left(k_{1}+k_{2}\psi\left(\displaystyle\frac{r}{\mu}\right)\right)\displaystyle\frac{\sqrt{\eta r^{2}+\epsilon\mu^{2}}}{r}+\epsilon+\displaystyle\frac{2\eta q^{2}}{\mu^{2}}-\Lambda r^{2}\right]$ (3.24)
$+2\eta k_{2}+4\epsilon\Lambda\mu^{2}.$
The possible singularities are at $r\to\infty$ and at the center $r=0$. As in case 1, $r\to\infty$ is a naked singularity at finite geodesic distance in subcases 2a or 2b unless $\Lambda=0$. If $\Lambda=0$ but $k_{2}\neq 0$, the logarithmic divergence of the function $\psi(r/\mu)$ (an arsinh or arcosh ) at $r\to\infty$ in subcases 2a or 2b carries over to the curvature invariants. However, the coordinate transformation $r=\mu e^{x}$ leads for $d=1/\mu^{2}$ to the asymptotic form of the metric
@@
showing that spacelike (for $k_{2}>0$) or timelike (for $k_{2}<0$) geodesics extend to $r\to\infty$. In other words, geodesics do not terminate at $r\to\infty$ if $\Lambda=0$ but $k_{2}\neq 0$. The invariants also diverge for $r\to 0$, unless $q=0$ and $k_{1}=0$.
@@ 3.2.1 Subcase 2a
We choose $\Lambda=0$ in (2.17) to avoid naked singularities for $r\to\infty$, and discuss first the simple case $k_{2}=0$. In the vacuum case $q=0$,
$y(r)=1+k_{1}\frac{\rho}{r}\quad\left(\rho=\sqrt{r^{2}+\mu^{2}}\right),$ (3.26)
with $k_{1}>-1$ for $y(\infty)>0$. For $-1<k_{1}<0$, the spacetime [2.15) is a black hole, with a single horizon shielding the singularity $r=0$. For $k_{1}>0$ the singularity is naked. For $k_{1}=0$, the singularity-free metric
@@
$y(r)=\bar{q}^{2}\frac{\rho^{2}}{r^{2}}+k_{1}\frac{\rho}{r}+\bar{q}^{2}+1\quad(\bar{q}=q/\mu),$ (3.28)
with $01+k_{1}+2\bar{q}^{2}>$ 0. For $-(1+2\bar{q}^{2})<k_{1}<-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$ the spacetime is a black hole with two horizons and a central singularity. It becomes an extreme black hole for $k_{1}=-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$, with a double horizon at $r=q$. For $k_{1}>-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$, the singularity is naked.
@@
where $y_{0}(r)$ is the function (3.26) or (3.28), and $z(r)=(\rho/r){\rm arsinh}\,(r/\mu)-1$ is a positive, steadily increasing function, such that $z(0)=z^{\prime}(0)=0$. It follows that the results obtained above for $k_{2}=0$ remain qualitatively valid (without a lower bound on $k_{1}$). For $q=0$ the spacetime is a black hole with a single horizon if $k_{1}<0$, it is geodesically complete if $k_{1}=0$, and the singularity is naked if $k_{1}>0$. For $q\neq 0$ the spacetime corresponds either to a black hole with a double horizon, to an extreme black hole, or to a naked singularity, depending on the parameter values.
@@ 3.2.2 Subcase 2b
We again choose $\Lambda=0$ to avoid naked singularities for $r\to\infty$. In this case, because $r^{2}>\mu^{2}$, the metric is best parameterized in terms of the radial coordinate $\rho=\sqrt{r^{2}-\mu^{2}}$:
@@
$y(\rho)=\bar{q}^{2}x^{2}+k_{1}x+\bar{q}^{2}-1+k_{2}[x\,{\rm artanh}\,{x}-1],\;\;x=\displaystyle\frac{\rho}{\sqrt{\mu^{2}+\rho^{2}}}\;\;(-1<x<1).$ (3.31)
@@
We first discuss the case $k_{2}=0$. Because of the quadratic form of the metric function $y$, there are at most two horizons. The central singularity is absent, so that the spacetime will be geodesically complete after analytical extension. The solution is stationary for $\rho\to+\infty$ provided
$k_{1}>1-2\bar{q}^{2},$ (3.32)
which we now assume. It will also be stationary for $\rho\to-\infty$ if $k_{1}<2\bar{q}^{2}-1$, in which case there will be two or no horizons. On the contrary, if $k_{1}>2\bar{q}^{2}-1$, the metric function $y(\rho)$ changes sign between $\rho\to+\infty$ and $\rho\to-\infty$, so that there is a single horizon.
We consider first the second possibility which, on account of (3.32), is realized if
$\bar{q}^{2}\leq\frac{1}{2}\quad{\rm or}\quad\bar{q}^{2}>\frac{1}{2}\;\;{\rm and}\;\;k_{1}>2\bar{q}^{2}-1.$ (3.33)
This case includes the vacuum solution $q=0$. The spacetime is a singularity-free black hole with one horizon.
In the complementary case
$\bar{q}^{2}>\frac{1}{2}\;\;{\rm and}\;\;k_{1}<2\bar{q}^{2}-1,$ (3.34)
there are three possibilities. For
$k_{1}^{2}>4\bar{q}^{2}(\bar{q}^{2}-1),$ (3.35)
the spacetime is a singularity-free black hole with two horizons. For
$k_{1}^{2}=4\bar{q}^{2}(\bar{q}^{2}-1)$ (3.36)
the spacetime is a singularity-free extreme black hole. And for
$k_{1}^{2}<4\bar{q}^{2}(\bar{q}^{2}-1),$ (3.37)
the spacetime is a traversable wormhole, which is symmetrical if $k_{1}=0$, the wormhole condition (3.37) implying $\bar{q}^{2}>1$. The existence of a traversable wormhole solution to conformal Killing gravity is a remarkable feature, since these are forbidden in general relativity by the null energy condition [23].
@@ 3.2.3 Subcase 2c
This case is complementary to the preceding, with $r^{2}<\mu^{2}$. The metric is, in terms of the radial coordinate $\psi=\arcsin{r/\mu}$ ($-\pi/2<\psi<\pi/2$),
@@
$y(\psi)=\bar{q}^{2}\cot^{2}\psi+k_{1}\cot\psi+1-\bar{q}^{2}+k_{2}(\psi\cot\psi-1)-\displaystyle\frac{\Lambda\mu^{2}}{3}\sin^{2}\psi.$ (3.41)
As the spacetime does not extend to spacelike infinity, there is no constraint on the integration constant $\Lambda$. If $q$ and/or $k_{1}$ are different from 00, the metric is singular at the center $\psi=0$. More interesting is the case of the vacuum solution, $q=0$, with also $k_{1}=0$. The metric is then singularity-free. It is stationary at the center, where $y(0)=1$, and is stationary everywhere ($y(\psi)>0$) for all $\psi\in[-\pi/2,\pi/2]$ provided
$k_{2}+\displaystyle\frac{\Lambda\mu^{2}}{3}<1.$ (3.42)
This case corresponds to a regular, closed static universe, with finite volume $V=2\pi^{2}\mu^{3}$. In the special subcase $\Lambda=k_{2}=0$, this is the Einstein static universe.

### v3

@@
Gérard Clémenta and Khireddine Nouicerb
a LAPTh, Université Savoie Mont Blanc, CNRS, F-74940 Annecy, France
b LPTh, Department of Physics, University of Jijel,
@@ 1 Introduction
In the pursuit of understanding the universe’s accelerated expansion [2, 3], numerous alternatives have been proposed, both within the framework of general relativity and beyond. One such alternative involves introducing a new matter component termed dark energy, characterized by negative pressure. While the cosmological constant $\Lambda$, originally introduced by Einstein, remains the primary candidate for dark energy due to its alignment with observational data, challenges such as the fine-tuning and coincidence problems persist [4, 5, 6, 7]. Recently, the DESI collaboration [8] has found evidence for dynamical dark energy in the form of quintessence. This finding implies that dark energy may not be a constant, necessitating a re-examination of the concordance model and offering new insights into the universe’s expansion and ultimate fate. On the other hand, a different path follows the idea that general relativity fails to describe the universe at larger scales and hence it should be modified. These modifications often propose alterations to fundamental principles of physics, seeking to reconcile observational evidence with theoretical predictions [9, 10]. Consequently, alternative models like scalar tensor theories [11, 12], Einstein-æther theory [13, 14], and higher derivative theories of gravity like $f(R)$ theories [15, 16] have been explored. Other kinds of modified theories are the modified teleparallel equivalent of general relativity, $f(T)$ gravity where the Ricci scalar is replaced by the torsion scalar, and the theory grounded in Weyl geometry, termed $f(Q)$ gravity [17]. The latter describes gravitational effects in terms of non-metricity, providing a broader framework than traditional Riemannian geometry. These theories offer different perspectives on gravity that may lead to new insights into cosmological and astrophysical phenomena.
@@ 2 Static spherically symmetric solutions with Maxwell source
Case 2: $d\neq 0,\,c\neq 0.$ The metric has the Lorentzian signature in three subcases: 2a) $c>0,\,d>0$, with $r$ real; 2b) $c<0,\,d>0$, with $r^{2}>-c/d$; 2c) $c>0,\,d<0$, with $r^{2}<-c/d$. In these three cases, the modulus of $c$ can be set to $1$ by a rescaling of the time coordinate, so we can fix $c=\epsilon=\pm 1$. Then, defining $\mu^{2}=1/|d|$, one may write $d=\eta/\mu^{2}$, where $\eta=\pm 1$, and $\mu$ is a constant with the dimension of a length. The generic solution for these three cases is
@@
$+k_{2}\left[\displaystyle\frac{\sqrt{\eta r^{2}+\epsilon\mu^{2}}}{r}\,\psi\left(\displaystyle\frac{r}{\mu}\right)-1\right]+k_{3}r^{2},$
@@
$a)\psi(x)=$ ${\rm arcsinh}\,{x},\quad\epsilon=+1,\;\eta=+1,$ (2.18)
$b)\psi(x)=$ ${\rm arccosh}\,{x},\quad\epsilon=-1,\;\eta=+1$ $\quad(r^{2}>\mu^{2}),$ (2.19)
@@
Expanding the solutions 2a and 2c with $\epsilon=+1$ in Laurent series for the variable $r/\mu$, we recover the solutions given by Barnes (equation (20) of [23]):
$y(r)=1+(1+2d\,r^{2})q^{2}/r^{2}-2mp_{1}(r)+\lambda p_{2}(r)r^{4}/5-\Lambda r^{2}/3,$ (2.21)
where $m=-k_{1}\mu/2$, $\lambda=-2k_{2}/3\mu^{4}$, $\Lambda=-\eta k_{2}/\mu^{2}-3k_{3}$, and the series $p_{1}(r)$ and $p_{2}(r)$ are summed up as
$p_{1}(r)$ $=$ $\sqrt{1+d\,r^{2}},$ (2.22)
$p_{2}(r)$ $=$ $-\displaystyle\frac{15}{2dr^{4}}\left[\displaystyle\frac{\mu\sqrt{1+d\,r^{2}}}{r}\psi\left(\displaystyle\frac{r}{\mu}\right)-1-\displaystyle\frac{d\,r^{2}}{3}\right].$ (2.23)
The solution 2b with $\epsilon=-1$, $\eta=+1$ was previously given by Barnes (equation (24) of [23]) for the special case $k_{1}=k_{2}=0$.
@@ 3 Geometry
The coordinate singularities of the metric (2.15) are at $r=0$ (possible curvature singularity), $y(r)=0$ (horizons), $y(r)\to\infty$ (possible curvature singularity), and $c+d\,r^{2}=0$ (cases 2b and 2c). We will see in the following that similarly to the case of regular horizons, where maximal analytic extension may be carried out by transforming to Kruskal-like coordinates, the metric (2.15) may be extended through the zeroes of $c+d\,r^{2}$.
@@ 3.1 Case 1
When $\lambda=0$ the charged Harada solution (2.16) reduces to the well-known Reissner-Nordström-(anti-)de Sitter solution of general relativity. It was suggested by Harada [18] that, when $\lambda\neq 0$, the new term in $r^{4}$ may become significant at large distances. However the Ricci scalar
@@
diverges on the sphere $r\to\infty$. Furthermore, the coordinate transformation $r=x^{-1}$ leads to the asymptotic form of the metric
@@
showing that this singular sphere is actually at finite geodesic distance for all geodesics if $\lambda>0$, and for spacelike radial geodesics if $\lambda<0$. If $m$, $\Lambda$ and $\lambda$ are negative, $y(r)$ is positive definite and the spacetime described by metric (2.15) ends on the two naked singularities $r=0$ and $r\to\infty$. In the generic case there can be up to six horizons between these two singularities. If we assume for simplicity $q=\Lambda=0$ and $m>0$, then for $\lambda<0$ the stationary domain is bounded on one side by a horizon shielding the Schwarzschild-like spacelike singularity $r=0$ and on the other side by the naked timelike singularity $r\to\infty$, while for $\lambda>0$ two horizons $r=r_{\pm}$ hide two spacelike singularities.
@@ 3.2 Case 2
$\mu^{2}R$ $=$ $-6\eta\left[\displaystyle\frac{\epsilon q^{2}}{r^{2}}+\left(k_{1}+k_{2}\psi\left(\displaystyle\frac{r}{\mu}\right)\right)\displaystyle\frac{\sqrt{\eta r^{2}+\epsilon\mu^{2}}}{r}+\epsilon+\displaystyle\frac{2\eta q^{2}}{\mu^{2}}+3k_{3}r^{2}\right]$ (3.3)
$+2\eta k_{2}-12\epsilon k_{3}\mu^{2}.$
The possible curvature singularities are at $r\to\infty$ and at the center $r=0$. Similarly to case 1, the singularity $r\to\infty$ is at finite geodesic distance in subcases 2a or 2b unless $k_{3}=0$. If $k_{3}=0$ but $k_{2}\neq 0$, the logarithmic divergence of the function $\psi(r/\mu)$ (an arcsinh or arccosh ) at $r\to\infty$ in subcases 2a or 2b carries over to the curvature invariants. However, the coordinate transformation $r=\mu e^{x}$ leads for $d=1/\mu^{2}$ to the asymptotic form of the metric
@@
showing that spacelike (for $k_{2}>0$) or timelike (for $k_{2}<0$) geodesics extend to $r\to\infty$. In other words, geodesics do not terminate at $r\to\infty$ if $k_{3}=0$ but $k_{2}\neq 0$. The invariants also diverge for $r\to 0$, unless $q=0$ and $k_{1}=0$.
@@ 3.2.1 Subcase 2a
We choose $k_{3}=0$ in (2.17) to avoid the curvature singularity at $r\to\infty$, and discuss first the simple case $k_{2}=0$. In the vacuum case $q=0$,
$y(r)=1+k_{1}\frac{\rho}{r}\quad\left(\rho=\sqrt{r^{2}+\mu^{2}}\right).$ (3.5)
For $k_{1}\leq-1$, there is no stationary domain ($y(r)$ is negative definite). For $-1<k_{1}<0$, the spacetime [2.15) is a non-asymptotically flat black hole, with a single horizon shielding the singularity $r=0$. For $k_{1}>0$ the singularity is naked. For $k_{1}=0$, the singularity-free metric
@@
$y(r)=\bar{q}^{2}\frac{\rho^{2}}{r^{2}}+k_{1}\frac{\rho}{r}+\bar{q}^{2}+1\quad(\bar{q}=q/\mu).$ (3.7)
For $k_{1}\leq-(1+2\bar{q}^{2})$, $y(\infty)<0$ so that the central timelike singularity is not shielded by the cosmological horizon. For $-(1+2\bar{q}^{2})<k_{1}<-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$ the spacetime is a black hole with two horizons and a central singularity. It becomes an extreme black hole for $k_{1}=-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$, with a double horizon at $r=|q|$. For $k_{1}>-2|\bar{q}|\sqrt{1+\bar{q}^{2}}$, the singularity is naked.
@@
where $y_{0}(r)$ is the function (3.5) or (3.7), and $z(r)=(\rho/r){\rm arcsinh}\,(r/\mu)-1$ is a positive, steadily increasing function, such that $z(0)=z^{\prime}(0)=0$. It follows that the results obtained above for $k_{2}=0$ remain qualitatively valid (without a lower bound on $k_{1}$). For $q=0$ the spacetime is a black hole with a single horizon if $k_{1}<0$, it is geodesically complete if $-1<k_{1}<0$, and the singularity is naked if $k_{1}>0$. For $q\neq 0$ the spacetime corresponds either to a black hole with a double horizon, to an extreme black hole, or to a naked singularity, depending on the parameter values.
@@ 3.2.2 Subcase 2b
We again choose $k_{3}=0$ to avoid the curvature singularity at $r\to\infty$. In this case, because $r^{2}>\mu^{2}$, the metric is best parameterized in terms of the radial coordinate $\rho=\sqrt{r^{2}-\mu^{2}}$,
@@
$y(\rho)=\bar{q}^{2}x^{2}+k_{1}x+\bar{q}^{2}-1+k_{2}[x\,{\rm arctanh}\,{x}-1],\;\;x=\displaystyle\frac{\rho}{\sqrt{\mu^{2}+\rho^{2}}}\;\;(-1<x<1).$ (3.10)
@@
We first discuss the case $k_{2}=0$. The central singularity is absent, so that the spacetime will be geodesically complete after analytical extension. Because of the quadratic form of the metric function $y$, there are at most two horizons. The extremal values of $y(\rho)$ are
$y(\pm\infty)=2\bar{q}^{2}-1\pm k_{1},\quad y_{\rm min}=\displaystyle\frac{4\bar{q}^{2}(\bar{q}^{2}-1)-k_{1}^{2}}{4\bar{q}^{2}}.$ (3.11)
From these values we see that the range of the parameter $|k_{1}|$ can be divided in three segments.
  1. 1.
    $k_{1}^{2}>(2\bar{q}^{2}-1)^{2}$. In this case, which includes the vacuum solution $q=0$, the spacetime is a singularity-free black hole with one horizon.
  2. 2.
    $4\bar{q}^{2}(\bar{q}^{2}-1)<k_{1}^{2}<(2\bar{q}^{2}-1)^{2}$ and $\bar{q}^{2}>1/2$ (for $\bar{q}^{2}<1/2$, $y(\rho)$ is negative definite so that there is no stationary domain). In this case, the spacetime is a singularity-free black hole with two horizons.
  3. 3.
    $k_{1}^{2}<4\bar{q}^{2}(\bar{q}^{2}-1)$. In this case, which occurs only if $\bar{q}^{2}>1$, $y(\rho)$ is positive definite so that the spacetime is a traversable wormhole, which is symmetrical if $k_{1}=0$.
The existence of a traversable wormhole solution to conformal Killing gravity is a remarkable feature, since these are forbidden in general relativity by the null energy condition [24]. To show explicitly that this wormhole is traversable, let us write down the first integrated geodesic equation for timelike equatorial ($\theta=\pi/2$) geodesics, with $E$ and $L$ the energy and angular momentum of a unit mass test particle, as
$\dot{\rho}^{2}+V(\rho)=0,$ (3.12)
where the effective potential is
$V(\rho)=\displaystyle\frac{\mu^{2}+\rho^{2}}{\mu^{2}}\left(y(\rho)-E^{2}\right)+\displaystyle\frac{L^{2}}{\mu^{2}}y(\rho).$ (3.13)
The metric function $y(\rho)$ is bounded by $y(\rho)\leq A$, where $A=2\bar{q}^{2}-1+|k_{1}|$ (the largest of the two values of $y(\rho)$ at infinity). It then follows that the first term of $V(\rho)$ is negative with the upper bound $A-E^{2}$ provided $E^{2}>A$, while the positive second term has the upper bound $L^{2}A/\mu^{2}$. So the effective potential $V(\rho)$ is negative definite in the whole real range of the radial coordinate $\rho$ if the energy is large enough, $E^{2}>A(1+L^{2}/\mu^{2})$.
@@ 3.2.3 Subcase 2c
This case is complementary to the preceding, with $r^{2}<\mu^{2}$. The metric is, in terms of the radial coordinate $\psi=\arcsin(r/\mu)$ ($20<\psi<\pi/$ 2),
@@
$y(\psi)=\bar{q}^{2}\cot^{2}\psi+k_{1}\cot\psi+1-\bar{q}^{2}+k_{2}(\psi\cot\psi-1)+\mu^{2}k_{3}\sin^{2}\psi.$ (3.17)
If $q$ and/or $k_{1}$ are different from 00, the metric has a curvature singularity at the center $\psi=0$. On the other hand, there is no curvature singularity at $\psi=\pi/2$, so that the range of $\psi$ may be extended (possibly with Kruskal-like analytical continuation through the zero(es) of $y(\psi)$) to $0<\psi<\pi$, $\psi=\pi$ being again a curvature singularity.
If $q=k_{1}=0$, the metric becomes regular at $\psi=0$, where $y(0)=1$. It follows that the range of $\psi$ may again be extended to $-\pi<\psi<\pi$, with again curvature singularities at the two ends $\psi=\pm\pi$ of this interval.
If furthermore $k_{2}=0$, the metric function reduces to
$y(\psi)=1-\displaystyle\frac{\Lambda\mu^{2}}{3}\,\sin^{2}\psi.$ (3.18)
The singularity-free spacetime is then geodesically complete (possibly after Kruskal continuation). It is stationary everywhere if $\Lambda<3/\mu^{2}$. This case corresponds to a regular, closed static universe, with finite volume $V=2\pi^{2}\mu^{3}$. In the special subcase $k_{3}=0$ ($\Lambda=0$), this is the Einstein static universe.